OpenStudy (anonymous):

When you solve the equation x plus nine equals fifteen, how do you know what number to subtract from both sides? Why is it important to subtract the same number from both sides of the equation?

OpenStudy (anonymous):

x + 9 = 15 x + 9 - 9 = 15 - 9 x = 6 You would subtract the nine from both side because you want to solve for x. And to solve for x you have to get x by itself.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

No problem. :) Hope that helped.